# Emulating Dark Matter Halo Merger Trees with Graph Generative Models

### v1

@@ Abstract
pubyear: 2025pagerange: Emulating Dark Matter Halo Merger Trees with Graph Generative ModelsEmulating Dark Matter Halo Merger Trees with Graph Generative Models
@@ 1 Introduction
There are two main approaches for generating DM-only merger trees that represent structure formation in a chosen $\Lambda$CDM cosmology: (1) extracting them cosmological DM-only (or $N$-body) simulations and (2) constructing them semi-analytically using various methods based on the Extended Press-Schechter (EPS) formalism (Press & Schechter, 1974; Bond et al., 1991; Bower, 1991).
@@
However, simulations have limited dynamic range due to computational costs that scale steeply with simulated volume and mass resolution. Large volumes improve sample statistics and are required for capturing rare, massive halos such as galaxy clusters ($10^{14}-10^{15}\,\mathrm{M_{\odot}}$), while high resolution is necessary to resolve halos at galactic scales. Due to the hierarchical nature of $\Lambda$CDM, accurate modeling of galaxy formation requires resolving halos down to masses of at least $\sim 10^{10}\,\mathrm{M_{\odot}}$. This challenge becomes even more pronounced at high redshifts, where halos are less massive and more abundant, requiring higher mass resolution and finer temporal sampling to accurately track their evolution. Consequently, simulations must balance resolution and volume, limiting their ability to simultaneously capture the full range of halo properties across cosmic time.
@@ 2.3 Training and test datasets
Similar to N24, we implement snapshot subsampling to reduce computational demands and help improve model generalization by preventing overfitting to specific snapshot sequences. For training, we randomly subsample snapshots at uniform intervals of $2-6$ snapshots, extending to $z=10$. When snapshots are removed, progenitor halos from retained snapshots are directly linked to descendants of their immediate descendants, maintaining causal structure. Subsampling is performed once per tree, though multiple iterations can increase data volume when needed (e.g., Appendix B). For testing, we subsample every 4 snapshots extending to $z=10$ (27 snapshots), to facilitate more straightforward comparisons. We discuss implications of this strategy in Section 5.
@@ 4.1 Progenitor mass distributions
For each tree, we extract all descendant-progenitor and progenitor-progenitor pairs and put them into bins according to their root mass $M_{\text{r}}$ and descendant redshift $z_{\text{d}}$. We use four mass bins, with edges (10,11, 12, 13, $(10\,,11,\,12,\,13,\,14)\,\mathrm{\mathrm{M_{\odot}}dex}$, and two redshift bins, with edges (0.00, 1.43, 4.88)0.001.434.88(0.00,\,1.43,\,4.88)( 0.00 , 1.43 , 4.88 ). Figures 3 and 4 show the distributions of $\mu_{i}$ and $\mu_{i,j}$, respectively for $i,j=1,2,3$.
@@ 4.2 Merger rates
Figure 5 compares the merger rates $B(M_{\text{d}},\xi_{i},z_{\text{p}},z_{\text{d}})$ of FLORAH-Tree, P08 EPS, and VSMDPL trees across three progenitor-descendant redshift slices and five descendant mass $M_{\text{d}}$ bins, with error bars representing Poisson uncertainties. When computing $B$, we sum over the contributions from $i=2$ (secondary) and $i=3$ (tertiary) progenitors. To improve bin statistics, we use overlapping mass bins with lower edges (10.0, 10.5, 11.0, 11.5, 12.0)10.010.511.011.512.0(10.0,\,10.5,\,11.0,\,11.5,\,12.0)( 10.0 , 10.5 , 11.0 , 11.5 , 12.0 ) $\mathrm{dex}$ and a width of $2.0$ $\mathrm{dex}$ in $\mathrm{M_{\odot}}$.
@@ 5.1 Limitations of the current model
As discussed in Section 2, we construct merger trees by subsampling simulation snapshots at intervals of $2-6$ snapshots. This subsampling reduces the temporal resolution for tracking galaxy formation processes. Depending on the application, one might need higher temporal resolution to capture finer details of merger timing and halo assembly. In future work, we aim to increase this temporal resolution further.
@@ 6 Conclusions
    Section 4.1 examines the progenitor-descendant (Figure 3) and progenitor-progenitor (Figure 4) mass ratio distributions at several redshifts and root halo mass ranges. FLORAH-Tree successfully models these distributions with some minor discrepancies primarily towards distribution tails. Notably, the model performs consistently well across the full halo mass spectrum, despite the significant imbalance in our training dataset, where root halos with masses $>10^{12}\,\mathrm{M_{\odot}}$ represent only about 0.7% of the total merger trees. Furthermore, FLORAH-Tree maintains accuracy at high redshifts without exhibiting the error accumulation typical of autoregressive models during sequential predictions.
@@ B.1 The GUREFT simulations and dataset
[…]
@@
We present results for the GUREFT-05 box, which contains the highest number of merger trees. We extract approximately 35,000 merger trees from the simulation and split them $80-20$ for training/validation and testing, with a further $80-20$ split between training and validation. To augment the training dataset, we create additional samples by randomly sub-sampling each tree at intervals of $2-6$ snapshots, generating three realizations for each original tree. This results in approximately 100,000 merger trees in the training and validation sets in total.
@@ B.2 Results
[…]

### v2

@@ Abstract
pubyear: 2025pagerange: Emulating Dark Matter Halo Merger Trees with Graph Generative ModelsReferences
@@ 1 Introduction
[…]
@@
[…]
@@ 2.3 Training and test datasets
Similar to N24, we implement snapshot subsampling to reduce computational demands and help improve model generalization by preventing overfitting to specific snapshot sequences. For training, we randomly subsample snapshots at uniform intervals of 262-6 snapshots, extending to $z=10$. When snapshots are removed, progenitor halos from retained snapshots are directly linked to descendants of their immediate descendants, maintaining causal structure. Subsampling is performed once per tree, though multiple iterations can increase data volume when needed (e.g., Appendix B). For testing, we subsample every 4 snapshots extending to $z=10$ (27 snapshots), to facilitate more straightforward comparisons. We discuss implications of this strategy in Section 5.
@@ 4.1 Progenitor mass distributions
For each tree, we extract all descendant-progenitor and progenitor-progenitor pairs and put them into bins according to their root mass $M_{\text{r}}$ and descendant redshift $z_{\text{d}}$. We use four mass bins, with edges $(10\,,11,\,12,\,13,\,14)\,\mathrm{\mathrm{M_{\odot}}dex}$, and two redshift bins, with edges (0.00, 1.43, 4.88)(0.00,\,1.43,\,4.88). Figures 3 and 4 show the distributions of $\mu_{i}$ and $\mu_{i,j}$, respectively for $i,j=1,2,3$.
@@ 4.2 Merger rates
Figure 5 compares the merger rates $B(M_{\text{d}},\xi_{i},z_{\text{p}},z_{\text{d}})$ of FLORAH-Tree, P08 EPS, and VSMDPL trees across three progenitor-descendant redshift slices and five descendant mass $M_{\text{d}}$ bins, with error bars representing Poisson uncertainties. When computing $B$, we sum over the contributions from $i=2$ (secondary) and $i=3$ (tertiary) progenitors. To improve bin statistics, we use overlapping mass bins with lower edges (10.0, 10.5, 11.0, 11.5, 12.0)(10.0,\,10.5,\,11.0,\,11.5,\,12.0) $\mathrm{dex}$ and a width of $2.0$ $\mathrm{dex}$ in $\mathrm{M_{\odot}}$.
@@ 5.1 Limitations of the current model
As discussed in Section 2, we construct merger trees by subsampling simulation snapshots at intervals of 262-6 snapshots. This subsampling reduces the temporal resolution for tracking galaxy formation processes. Depending on the application, one might need higher temporal resolution to capture finer details of merger timing and halo assembly. In future work, we aim to increase this temporal resolution further.
@@ 6 Conclusions
    Section 4.1 examines the progenitor-descendant (Figure 3) and progenitor-progenitor (Figure 4) mass ratio distributions at several redshifts and root halo mass ranges. FLORAH-Tree successfully models these distributions with some minor discrepancies primarily towards distribution tails. Notably, the model performs consistently well across the full halo mass spectrum, despite the significant imbalance in our training dataset, where root halos with masses $10^{12}\,\mathrm{M_{\odot}}$ represent only about 0.7% of the total merger trees. Furthermore, FLORAH-Tree maintains accuracy at high redshifts without exhibiting the error accumulation typical of autoregressive models during sequential predictions.
@@ B.1 The GUREFT simulations and dataset
The GUREFT (Gadget at Ultrahigh Redshift with Extra-Fine Timesteps; Yung et al. 2024) simulations are a suite of $N$-body simulations specifically designed to probe the assembly histories of DM halos in the ultra-high redshift universe. The suite consists of four simulations with box sizes of $(5,\,15,\,35,\,90)\;\mathrm{Mpc}\,h^{-1}$ respectively [^7] This corresponds to $(7.37,\,22.1,\,51.6,\,133)\;\mathrm{Mpc}$ in physical units., each containing $1024^{3}$ particles with mass resolutions of $(6.7\times 10^{3},\,1.8\times 10^{5},\,2.3\times 10^{6},\,3.9\times 10^{7})\;\mathrm{M_{\odot}}$. A distinctive feature of GUREFT is its unprecedented temporal resolution, with the simulations storing 171 snapshots between $z\approx 6-40$, spaced at one-tenth of the halo dynamical time. This enables detailed reconstruction of merger histories during the earliest epochs of cosmic structure formation.
@@
We present results for the GUREFT-05 box, which contains the highest number of merger trees. We extract approximately 35,000 merger trees from the simulation and split them 802080-20 for training/validation and testing, with a further 802080-20 split between training and validation. To augment the training dataset, we create additional samples by randomly sub-sampling each tree at intervals of 262-6 snapshots, generating three realizations for each original tree. This results in approximately 100,000 merger trees in the training and validation sets in total.
@@ B.2 Results
We first examine the distributions of the descendant-progenitor mass ratio $\mu_{i}$ and the progenitor-progenitor mass ratio $\mu_{i,j}$ for i=1, 2, 3i=1,\,2,\,3 and $i>j$. Figures 7 and 8 presents these distributions, respectively, and compares the FLORAH-Tree results with GUREFT-05. Note that certain distributions with insufficient statistics (fewer than 100 samples) are omitted from the figures. Qualitatively, the performance of FLORAH-Tree is similar to that observed with the VSMDPL dataset, where the overall distributions are well-reproduced but with some discrepancies towards the tails of the $\mu_{i}$ distributions. We again do not observe significant differences in the performance of FLORAH-Tree across different mass and redshift bins.